OpenStudy (anonymous):

i got the answer to this question wrong A scientist counts 35 bacteria present in a culture and finds that the number of bacteria triples each hour. The function y = 35 • 3x models the number of bacteria after x hours. Graph the function. (You do not need to submit the graph). Use the graph to estimate when there will be about 550 bacteria in the culture.

OpenStudy (anonymous):

it was my answer y= 35 . 3^x (35 is the initial number of bacteria) y= 35 . 3^(1) (substitute when x=1, so that u can graph it) y=105 when x=1 (these r the points which u r goin to plot in the graph) then y=35.3^x (substite when x=2) y = 35 * 3^2 y = 35 * 9 = 315

OpenStudy (anonymous):

whats wrong with my answer

hartnn (hartnn):

you need to estimate when there will be about 550 bacteria in the culture. so, put y= 550 550 = 35. 3^x

hartnn (hartnn):

you first drew the graph, right ?

hartnn (hartnn):

those 2 points you calculated are correct.

OpenStudy (anonymous):

that would be 15.58

hartnn (hartnn):

15.58 for what ? x ? no....

OpenStudy (anonymous):

my teacher commented this You are trying to use the equation to estimate. You need to use the graph. I would suggest using a graphing utility. Find the desired y-value on the graph and determine what the corresponding x-value is.

OpenStudy (anonymous):

for x=15.58

hartnn (hartnn):

yeah, thats why i asked whether you drew the graph ?you

OpenStudy (anonymous):

i did but iam getting confused

hartnn (hartnn):

no, not for x=15.58, you see y=550 ON GRAPH, and find corresponding x co-ordinate

hartnn (hartnn):

for that you need graph first.

OpenStudy (anonymous):

can you tell me the equation s first

hartnn (hartnn):

there are no equations.... you need to graph y= 35 * 3^x using a graphing utility. then locate y= 550 point on that. do you have any graphing utility with you ?

OpenStudy (anonymous):

no

hartnn (hartnn):

neither do i have. when you have one, try this problem again.